Common stock value Variable growth   Lawrence​ Industries' most recent annual dividend was ​$1.54 per share ​(D0=$ 1.54​), and the​ firm's required return is 10​%. Find the market value of​ Lawrence's shares when dividends are expected to grow at 20​% annually for 3​ years, followed by a 7​% constant annual growth rate in years 4 to infinity.

The market value of​ Lawrence's shares is ​$
Answer #1

Do = 1.54 De = 1.54 x1.20 = 1.85 D = 1.85 x 1.20 -2.22 Dz = 2-22 x1.20 = 2.66 Dy = 2.66X1.07 - 2.85 Terminal volue (13) = Re